OpenStudy (anonymous):

Please perform the indicated operation AND state the domain. 4x^(2/3)/5x^(1/2) Thanks. Please perform the indicated operation AND state the domain. 4x^(2/3)/5x^(1/2) Thanks. @Mathematics

OpenStudy (mertsj):

Multiply the coefficients: 4 x 5 is 20 multiply the variables by adding the exponents: x^2/3(x^1/2) =x^7/6. So the result is 20x^7/6. Now the domain: since you have x^1/2, you are taking an even root of x. In the set of real numbers you cannot take the square root of a negative number so the domain is x must be greater than or equal to 0.

OpenStudy (anonymous):

Its actually division--sorry if that was unclear. My answer was 4x^(1/6)/5 The thing i dont understand is domain.

OpenStudy (mertsj):

What is the square root of -4 greeneyes?

OpenStudy (anonymous):

This has nothing to do with square roots...

OpenStudy (mertsj):

Yes it does sweetheart because x^1/2 means the square root of x

OpenStudy (anonymous):

(4/5)x^(1/6) domain = (0, infinity) because negative x=imaginary numbers

OpenStudy (anonymous):

the simplification of 4x^(2/3) / 5x^(1/2) is simply (4/5) * x^( (2/3) - (1/2) )

myininaya (myininaya):

\[\frac{4x^\frac{2}{3}}{5x^\frac{1}{2}}\] domain is x>0 --------------------------------- \[\frac{4}{5} \cdot \frac{x^\frac{2}{3}}{x^\frac{1}{2}}=\frac{4}{5} \cdot x^{\frac{2}{3}-\frac{1}{2}}=\frac{4}{5}x^{\frac{4}{6}-\frac{3}{6}}=\frac{4}{5}x^\frac{1}{6}\]

myininaya (myininaya):

mertsj is right \[\sqrt{x}=x^\frac{1}{2}\]

OpenStudy (anonymous):

I know the answer is 4x^(1/6)/5 the domain is all positive real #'s?

OpenStudy (anonymous):

yea it's 0 inclusive. so it should actually be [0, infinity) (note the usage of bracket and not parenthesis)

myininaya (myininaya):

\[\sqrt[a]{x^b}=x^\frac{b}{a}\] yes domain is all positive real numbers it doesn't include 0

OpenStudy (mertsj):

NO. The domain is x > 0 because you cannot have the square root of a negative number and x^1/2 means the square root of x so x cannot be negative and it cannot be 0 because it is in the denominator

myininaya (myininaya):

\[\frac{4x^\frac{2}{3}}{5 x^\frac{1}2}\] is the expression we are looking at

OpenStudy (mertsj):

Yes and if x = 0 then the denominator is 0 and the expression is meaningless.

myininaya (myininaya):

this expression equals \[\frac{4}{5}x^\frac{1}{6}\] as long as x does not equal 0

OpenStudy (mertsj):

Or as long as x is not negative

OpenStudy (mertsj):

You cannot have an even root of a negative root in the set of real numbers

OpenStudy (mertsj):

I mean negative number

OpenStudy (anonymous):

0 is inclusive because the x value is not in the denominator...

OpenStudy (anonymous):

hey i feel on simplifying the expression we are having \[0.8 \times \sqrt[6]{x}\] and as x is a variable so global minimum will be Zero

myininaya (myininaya):

i know they have the same domain if x=0 is not in question

OpenStudy (mertsj):

It is in the original expression. You cannot take something which is meaningless and make it meaningful by computing with it.

OpenStudy (mertsj):

Franklin. x cannot be 0

OpenStudy (anonymous):

x can be zero @ mertsj

myininaya (myininaya):

so fronk you are saying \[\frac{4}{5} \cdot \frac{0^\frac{2}{3}}{0^\frac{1}{2}} \text{ is a number that xeists}\]

myininaya (myininaya):

x cannot be zero

OpenStudy (mertsj):

It cannot.

OpenStudy (anonymous):

why dont u guys simplify the whole expression

myininaya (myininaya):

we did but you must consider the domain of the original expression

OpenStudy (mertsj):

We have done that numerous times. How many times do you require?

myininaya (myininaya):

(and yes mertsj I know they that the domains of both do not include the negatives) the two expression are = as long as we don't include x=0

OpenStudy (anonymous):

see if u are going to simplify then also u will be having only original expressions value in the same proportion as it is

OpenStudy (anonymous):

Well thanks for helping :P

OpenStudy (anonymous):

I know this is an old question, and i'm just asking b/c i need to know by tomorrow and people dont usually answer me...So, can one of u help me find the domain for 5x^(2/3) and -9x^(2/3)? Thanks :)